# A mathematical model for wind-generated particle-fluid flow fields with an application to the helicopter cloud problem

### v1

@@ 6 The helicopter cloud problem
7 Numerical method and examples
@@
[…]
@@
In our first example, we choose the dimensionless parameters as follows:
@@
[…]
@@
As $N$ increases, we see the error decrease, but the rate of convergence is somewhat erratic. To see more clearly how the solution converges, in figure 8 we plot the solution for $z_{d}=0.5$, for $N=2$, $N=4$, $N=8$ and $N=16$. These are plotted on the computational domain, which grows with increasing $N$. To ease comparison, we additionally plot (on the same figure) the $N=8$ and $N=16$ solutions over the same range as for the $N=4$ solution. In figure 9 we plot the relative errors $|(E_{N}-E_{16})/E_{16}|$, each over the same range as for the $N=2$ solution. We repeat the same plots for $z_{d}=0.9$ in figures 10-11, and for $z_{d}=1.3$ in figures 12-13.
Figure 8: Numerical solution $E_{N}$, $z_{d}=0.5$, example 1, computed for $N=2$ (top left), $N=4$ (top right), $N=8$ (middle left) and $N=16$ (middle right), each shown on the full computational domain. The $N=8$ (bottom left) and $N=16$ (bottom right) solutions are also plotted over the same range as for the $N=4$ solution, for easier comparison.
Figure 9: Relative error $|(E_{N}-E_{16})/E_{16}|$, for $z_{d}=0.5$, example 1, computed for $N=2$ (top left), $N=4$ (top right), $N=8$ (bottom). Note the different scales on the color bar for each plot.
@@
Figure 12: Numerical solution $E_{N}$, $z_{d}=1.3$, computed for $N=2$ (top left), $N=4$ (top right), $N=8$ (middle left) and $N=16$ (middle right), example 1, each shown on the full computational domain. The $N=8$ (bottom left) and $N=16$ (bottom right) solutions are also plotted over the same range as for the $N=4$ solution, for easier comparison.
Figure 13: Relative error $|(E_{N}-E_{16})/E_{16}|$, for $z_{d}=1.3$, example 1, computed for $N=2$ (top left), $N=4$ (top right), $N=8$ (bottom). Note the different scales on the color bar for each plot.
For each value of $z_{d}$, we see broadly comparable results. In each case, for $N=2$ the maximum relative error (as calculated in table 1) is where $R$ is largest and $z$ is smallest (lower right corner of plot) - this is because the domain is not large enough in this case to adequately capture the complete boundary data. As we increase $N$, the maximum relative error moves closer to the location of the helicopter rotor. The norm we have used in table 1 to measure error, namely the maximum pointwise relative error, solely measures the relative error at the worst point in the domain. To present a more balanced picture of the effectiveness of our method, illustrating numerically what we have seen qualitatively in figures 9, 11 and 13, in table 2 we focus on three specific points in the domain, to see how our numerical solution converges at those points. The values of $h$, DOF and iterations are the same in table 2 and 1.
[…]
@@
[…]
@@
Having established convergence of our numerical scheme in example 1, we now consider how the solution behaviour depends on our parameter choices. In examples 24, we fix
$E_{s}=0.01,\quad\alpha=0.1,\quad\Delta=0.04,$
and consider the nine possible combinations of
$\gamma=1000,500,100,\quad\Omega=1,0.5,0.1$
(note that $\gamma=500$, $\Omega=0.5$ corresponds to example 1). From our previous assessment of the numerical scheme, we consider it sufficient to choose the discretisation parameter $N=8$, and in order to ensure that the computational domain is large enough that all relevant solution behaviour will be captured across this wide range of parameter choices, we choose $L_{R}=32N$ and $L_{z}=16N$ (compared to $L_{R}=12N$ and $L_{z}=6N$ in example 1). These parameter choices lead to the total number of degrees of freedom in our numerical solution being $\mbox{DOF}=2317056$ for $z_{d}=0.5$, $\mbox{DOF}=2481536$ for $z_{d}=0.9$ and $\mbox{DOF}=2646016$ for $z_{d}=1.3$, for each of examples 24. As for example 1, we choose $\epsilon=1/(10N^{2})$. In each of the figures below, we truncate the computational domains where appropriate for presentational purposes, noting that the solution satisfies $E\approx 1$ everywhere outside the plotted range - note that all plots are over the same range within each figure, but that the plotted range varies between examples 24 ($R\in[0,6]$, $z\in[0,5]$ for example 2; $R\in[0,5]$, $z\in[0,4]$ for example 3; $R\in[0,3]$, $z\in[0,2]$ for example 4), though the computational domain for any given value of $z_{d}$ is identical for each example.
For all parameter choices in examples 2 and 3, and all in examples 4 except one (detailed below) we are either in case (a) or case (b)(i). This is illustrated for the case $\gamma=500$, $\Omega=0.5$, and $z_{d}=0.9$ or $z_{d}=0.5$ (as in example 1) in figure 14: for $z_{d}=0.9$, $\Omega=0.5>\Omega_{c}(0.9)$, and hence case (a) holds; for $z_{d}=0.5$, $\Omega=0.5<\Omega_{c}(0.5)$, and hence case (b) holds, however in this case both the local maximum and the local minimum take values greater than $(\gamma E_{s})^{-1}$ and $\gamma^{-1}$, hence case (b)(i) holds. The boundary data $g$ (shown in the lower part of figure 14) is qualitatively comparable in each of these cases.
Example 2 (Large $\gamma$)
Example 3 (Medium $\gamma$)
Example 4 (Small $\gamma$)
Again, we see more sand in the air (lower values of $E$) corresponding to larger values of $\Omega$, and to lower values of $z_{d}$. Comparing to figures 15 and 16, we see less sand in the air for the lower value of $\gamma$. Note that the plotted range is smaller in figure 17 than for both figure 16 and figure 15, though the computational domain is identical in each case. As for examples 2 and 3, for $\Omega=1$, and for $\Omega=0.5$ and $z_{d}=0.9,1.3$, the boundary data fits case (a), whilst for $\Omega=0.5$ and $z_{d}=0.5$, and for $\Omega=0.1$ and $z_{d}=0.9,1.3$ the boundary data fits case (b)(i), with similar qualitative solution behaviour. However, for $\Omega=0.1$ and $z_{d}=0.5$ we are now in case (b)(iii), as illustrated in figure 18. In this case, we see that the local maximum is greater than $(\gamma E_{s})^{-1}$, whilst the local minimum is less than $(\gamma E_{s})^{-1}$. As a result, the boundary data $g$ has a small hump near $X\approx 0.05$, though it is hard to see any discernible effect from this on the qualitative solution behaviour (bottom right plot of figure 17).
Example 5 (Two-plume example)
In our final example, we choose
$E_{s}=0.01,\quad\Delta=0.04,\quad\gamma=70,\quad\Omega=0.1,$
Figure 20: Voidage field $E$, example 5, $\alpha=0.1$, $z_{d}\in[0.4,1.1]$.
Figure 21: Voidage field $E$, example 5, $\alpha=0.02$, $z_{d}\in[0.4,1.1]$.
8 Conclusion
To illustrate the applicability of this theory we have formulated it in relation to a simple model of the helicopter cloud problem, leading to a fully determined nonlinear elliptic boundary value problem for the voidage field throughout the fluidized region. This provides a rational and efficient approach to computing the voidage field for this application. Careful numerical solution via a finite difference approximation scheme has demonstrated through a series of experiments how the qualitative behaviour of the voidage field varies as key parameters change, in a way that is entirely consistent with our understanding of the underlying physical processes. It may be of interest to look at this example in relation to controlled scale model experiments in order to specifically determine the qualitative and quantitative accuracy of the rational model that has been developed in this paper, but such considerations are left to future work.
[…]

### v2

@@ 8 Conclusion
[…]
[Data availability statement]The authors confirm that the data supporting the findings of this study are available within the article.
[…]
Appendix A The problem [BVP] in cylindrical polar coordinates $(R,\theta,z)$
On adopting (104), the problem [BVP] becomes, in terms of the cylindrical polar coordinates $(R,\theta,z)$,
$E=g(R)\mbox{ on }z=0,\quad R\geq 0,$ (114)
$E_{z}=0\mbox{ on }z=z_{d}^{+},z_{d}^{-},\quad R\in(0,\Delta),$ (118)
Here
for $(R,z)\in\omega$, whilst
$z_{d}^{+}=z_{d}+\frac{\Delta}{2},\quad z_{d}^{-}=z_{d}-\frac{\Delta}{2},$
and
[…]
Appendix B The numerical method, convergence and error estimates
@@
If we choose $L_{R},L_{z}\propto N$, then, recalling (101), we truncate at $R_{\infty},z_{\infty}=\mbox{{O}}(h^{-1})$, and in (127) take $\epsilon=\mbox{{O}}(h^{2})$. With mesh size $h$, and each step of our finite difference approximation accurate to O$(h^{2})$, we thus anticipate that as we decrease $h$ a reasonable expectation would be that the overall error in our numerical scheme is O$(h^{2})$. We test this in the following numerical example.
@@
In this example, we choose the dimensionless parameters as follows:
@@
$\mbox{EOC}=\log_{2}\left|\frac{\|(E_{N}-E_{16})/E_{16}\|_{\infty}}{\|(E_{2N}-E_{16})/E_{16}\|_{\infty}}\right|,$
@@
As $N$ increases, we see the error decrease, but the rate of convergence is somewhat erratic. To see more clearly how the solution converges, in figure 16 we plot the solution for $z_{d}=0.9$, for $N=2$, $N=4$, $N=8$ and $N=16$. These are plotted on the computational domain, which grows with increasing $N$. To ease comparison, we additionally plot (on the same figure) the $N=8$ and $N=16$ solutions over the same range as the $N=4$ solution. In figure 17 we plot the relative errors $|(E_{N}-E_{16})/E_{16}|$, each over the same range as the $N=2$ solution.
@@
For each value of $z_{d}$, we see broadly comparable results. In each case, for $N=2$ the maximum relative error (as calculated in table 1) is where $R$ is largest and $z$ is smallest (lower right corner of plot) - this is because the domain is not large enough in this case to adequately capture the complete boundary data. As we increase $N$, the maximum relative error moves closer to the location of the helicopter rotor. The norm we have used in table 1 to measure error, namely the maximum pointwise relative error, solely measures the relative error at the worst point in the domain. To present a more balanced picture of the effectiveness of our method, illustrating numerically what we have seen qualitatively in figure 17, in table 2 we focus on three specific points in the domain, to see how our numerical solution converges at those points. The values of $h$, DOF and iterations are the same in table 2 and 1.
$z_{d}$ $N$ $\mbox{error}_{1}$ EOC $\mbox{error}_{2}$ EOC $\mbox{error}_{3}$ EOC
@@
Table 2: Convergence of our numerical approximation scheme at three fixed points as $N$ increases, example 5. Here, $\mbox{error}_{m}=|(E_{N}(\mathbf{x}_{m})-E_{16}(\mathbf{x}_{m}))/E_{16}(\mathbf{x}_{m})|$, where $\mathbf{x}_{1}=(0.2,0.8)$, $\mathbf{x}_{2}=(0.5,0.5)$, $\mathbf{x}_{3}=(0.8,0.2)$.